### v3

# 142432, Chernogolovka, Moscow region, Russia Аннотация

Chiral twist-bend liquid crystals.
E.I.Kats Laue-Langevin Institute, F-38042, Grenoble, France, and
Landau Institute for Theoretical Physics, RAS,
142432, Chernogolovka, Moscow region, Russia
Аннотация
Recently published preprint (A. Ashkinazi, H.Chhabra, A.El.Moumane, M.M.C.Tortora, J.P.K. Doye, ’’Chirality transfer in lyotropic twist-bend nematics’’, arXiv:2508.03544v1 (2025)) has reawakened also interest to various mechanisms of chirality transfer from microscopic (molecular) level into the chirality of the structure. In this communication we present a simple theoretical analysis how the transfer occurs for the Landau model of phase transition between cholesteric ($Ch$) and chiral twist-bend ($N_{TB}^{*}$) liquid crystals. We found that the sign of the chiral heliconical $N_{TB}^{*}$ spiral is always opposite to that of the $Ch$. Physics behind this relation is based on the orthogonality of the cholesteric director and vector order parameter of the $N_{TB}^{*}$ phase.
Introduction.  The relatively recently discovered new type of liquid crystals with a tilted nematic-like structure (also known as twist-bend nematics) continue to attract the attention of researchers and continue to demonstrate new features of interest both for fundamental physics and for potentially promising applications HS09 ; PN10 ; CD11 ; BK13 ; MD14 . Especially surprising is the fact that the observed new phase structure demonstrate heliconical orientational order (therefore chiral) although they formed from achiral molecules. For comparison, well known conventional nematic liquid crystals ($N$) exhibit uniaxial and achiral orientational order, and also known for over 100 years chiral cholesteric liquid crystals possess simple (orthogonal) helical structures with pitches in a few $\mu m$ range. The cholesteric structure appears as a result of relatively weak molecular chirality (that is why it has a relatively large pitch), and the swirl direction of the spiral (left or right) is determined by the sign of the molecular chirality. Unlike this situation, the $N_{TB}$ nematics are formed as a result of spontaneous chirality breaking, they have nanoscale pitches.
Quite recently, AC25 $N_{TB}$ liquid crystals were also discovered in lyotropic liquid crystals. The phase transition studied in this work is athermal, i.e. it is controlled not by temperature, but by the volume fraction of mesogenic (’’banana-like’’, anisotropic) molecules dissolved in an isotropic solvent. Moreover, the lyotropic nature of a liquid crystal also allows studying mixtures of chiral and achiral mesogens. Thus, in the work AC25 , the chiral $N_{TB}^{*}$ phase of a liquid crystal was discovered and identified. In this case the chirality is already broken in the locally nematic-like (i.e., twisted nematic or cholesteric $Ch$) phase. Therefore, domains of both signs of twist (clockwise or counter-clockwise) of the heliconical structure in the chiral $N_{TB}$ appear at the phase transition $Ch$ - $N_{TB}^{*}$ with non-equal probabilities. Another interesting observation made in the work AC25 is that the sign of chirality of the chiral $N_{TB}$ structure is opposite to the sign of chirality of the cholesteric, from which the chiral lyotropic $N_{TB}^{*}$ liquid crystal arises.
The discovered in AC25 lyotropic $N_{TB}^{*}$ liquid crystal raises the natural question on the possibility of the existence of thermotropic $N_{TB}^{*}$ liquid crystals as well. In what follows in our work, we formulate Landau-like theory describing $Ch$ - $N_{TB}^{*}$ phase transition. The formulated model is a generalization to the chiral case, Landau’s theory previously developed KL14 formulated for the transition from an achiral nematic $N$ to a $N_{TB}$ liquid crystal. We will show below that our theory predicts that the chirality signs are opposite in the cholesteric $Ch$ and in the chiral $N_{TB}^{*}$ structure (the same relationship as it was found in AC25 ).
A conceptually close problem, on the mutual interplay of two competing types of chirality at incommensurate space scales, was solved in the works KA96 - KL00 . Technically significantly more advanced examples considered in these papers include chiral hexatic columnar phases formed by long polymer molecules. The matter is that in the examples considered in these works, the consideration is complicated by the fact that both types of chirality are frustrated, that is, they can be compatible due to the nucleation of topological defects. On the contrary for $Ch$ - $N_{TB}^{*}$ phase transition, the chirality of the both structures can be compatible without topological defects.
Landau theory model for the $Ch$ - $N_{TB}^{*}$ phase transition.
We start with a system (liquid crystal) in the cholesteric phase, where director ${\bf n}\to n_{x}=sin(q_{0}z)\,,n_{y}=cos(q_{0}z)\,,n_{z}=0)$ provides soft long-wavelength degrees of freedom. The $N_{TB}^{*}$ phase modulation is short-wavelength (i.e., on the order of molecular scale $a$), with its modulation wave vector $q_{s}a\simeq 1$ (in a contrast to the long-wavelength modulation of the $Ch$ phase, where $q_{0}a\ll 1$).
Since the modulation in the $N_{TB}^{*}$ structure is short wavelength, the corresponding order parameter is also the short wavelength one, and the free energy expansion should include the following terms KL14 :
  • Conventional, long wavelength Frank energy
    ${\cal F}_{\mathrm{Fr}}=\int dV\ \left\{\frac{K_{1}}{2}(\nabla\cdot\bm{n})^{2}+\frac{K_{2}}{2}\left[{\bm{n}}curl{\bm{n}}\right]^{2}+\frac{K_{3}}{2}(\partial_{z}\bm{n})^{2}\right\}.$ (1)
[…]
  • Non-gradient terms of the Landau free energy expansion for short wavelength order parameter have the same universal form as for usual long wavelength order parameter
    $F_{0}=\int dV\left\{\frac{a}{2}\bm{\varphi}^{2}+\frac{\lambda}{24}{\bm{\varphi}}^{4}\right\},$ (2)
    as it was said above since $\bm{\varphi}$} is a vector, odd order terms are forbidden.
  • On the contrary, the gradient terms for the short wavelength order parameter are different from those for the long wavelength order parameter. We assume that the $N_{TB}^{*}$ structure modulation wave vector oriented perpendicular to the local cholesteric director. Then the main term describing the softening of the order parameter in the vicinity of two points ${\bf q}=\pm q_{s}{\hat{e}}_{z}$ (where ${\hat{e}}_{z}$ is the unit vector along $Z$ axis chosen along the non-perturbed cholesteric axis) reads as
    $\int dV\left\{\frac{b_{||}}{8q_{s}^{2}}\left[\left(n_{i}n_{k}\partial_{i}\partial_{k}+q_{s}^{2}\right)\bm{\varphi}\right]^{2}\right\}.$ (3)
    However because for the short wavelength order parameter $\bm{\varphi}$} its gradients are not small, the contributions (2) should be supplemented by the following 2-d order terms
    $\int dV\left\{\frac{b_{1}}{2}(\nabla\bm{\varphi})^{2}+\frac{b_{2}}{2}\delta^{\perp}_{ij}\partial_{i}\bm{\varphi}\partial_{j}\bm{\varphi}\right\},$ (4)
    where $\delta^{\perp}_{ij}=\delta_{ij}-n_{i}n_{j}$. The terms (4) are the same order as the quadratic term $\propto a{\bm{\varphi}}^{2}$. Similarly the 4-th order term
    $\int dV\left\{\frac{\lambda_{1}}{16q_{s}^{2}}\left(\epsilon_{ijk}\varphi_{i}\partial_{j}\varphi_{k}\right)^{2}\right\},\quad$ (5)
    is generally of the same order as the term $(\lambda/24){\bm{\varphi}}^{4}$.
All enumerated above contributions presented above (1), (2 - 5) are achiral. To describe $Ch$ - $N_{TB}^{*}$ phase transitions, these terms should be supplemented by the chiral parts of the free energy. They read as
${\cal F}_{\mathrm{Chiral}}=\int dV\ \left\{\frac{K_{2}}{2}q_{0}\left[{\bm{n}}curl{\bm{n}}\right]+\frac{\lambda_{1}}{16q_{s}}\varepsilon_{ijk}\varphi_{i}\nabla_{j}\varphi_{k}\right\}.$ (6)
Having the above formulas in hands, we must carry out the minimization of the total free energy with respect to all admissible variations of the cholesteric director $\bm{n}$} and the vector order parameter $\bm{\varphi}$} for the $N_{TB}^{*}$ phase. However, since our aim in this work is very modest (to find the relationship between the chirality signs of the local cholesteric director $\bf n$} and the vector order parameter $\bm{\varphi}$}) we can avoid the explicit solutions of the corresponding Euler - Lagrange equations. To determine relation between the rotations signs for $\bm{n}$} and $\bm{\varphi}$}, we will exploit in the next section only a constraint between these quantities satisfying the Euler - Lagrange equations.
Chirality transfer from $Ch$ into $N_{TB}^{*}$ phase.
A direct, but rather cumbersome, minimization of the free energy presented above allows us to reach the following conclusion. At the phase transition from a cholesteric to a chiral $N_{TB}^{*}$ structure, the energy of the domains of the heliconical structure twisted clockwise or counterclockwise have different energies (in contrast to the $N$ - $N_{TB}$ phase transition, where these domains are degenerate in energy, as expected for spontaneous symmetry breaking). Moreover, for a cholesteric clockwise twist ($q_{0}>0$), the $N_{TB}^{*}$ domain twisted in the opposite direction has a lower energy. Similarly, for a cholesteric counterclockwise twist, domains of the $N_{TB}^{*}$ phase twisted in the clockwise direction has a lower energy. In order not to lose sight of physical results in algebraic exercises, we can rationalize the same observation using the simplest way to find the chirality sign of the vector order parameter $\bm{\varphi}$} describing the ordering in the chiral $N_{TB}^{*}$ phase. Namely by simple inspection of the complete free energy one can find the following conditions for the optimal structure of the $N_{TB}^{*}$ phase
$({\bm{n}}{\bm{\varphi}})=0,\quad$ (7)
Then, for a simple cholesteric spiral
${\bf n}=\left\{n_{x}=sin(q_{0}z)\,,\,n_{y}=cos(q_{0}z)\,,\,n_{z}=0\right\}$
we get the following condition
$sin(q_{0}z-q_{s}z)=0,\quad$ (8)
To satisfy the condition (8)
$q_{s}-q_{0}=\pi N,\quad$ (9)
where $N$ is an integer (odd or even) number.
Let us assume that $q_{0}>0$, i.e. we are dealing with a right-handed (clockwise) cholesteric helix. Therefore, for any point along the cholesteric rotation axis ($Z$ axis), it follows from (9) that
$q_{s}z>\pi N\quad.$ (10)
The above relation show that for even $N$, the final (at the point $z$) orientation of $\bm{\varphi}$} lies in the 4-th quadrant of the plane, whereas for the odd $N$, the final orientation is in the 3-d quadrant. However in the both cases, the direction of the $\bm{\varphi}$} rotations is anticlockwise. Similar trivial arguments for the anticlockwise cholesteric $q_{0}<0$, $\bm{\varphi}$} rotation is clockwise, i.e., $q_{s}>0$. Note to the point that for weakly chiral microscopic building blocks of the cholesterics, the macroscopic phase handedness is expected ST76 to have the opposite handedness to the particles. On the contrary for strong molecular chirality the phase chirality is the same as that for the particles ST76 .
Conclusion.  To place our manuscript with respect to other publications in this field we have to mention again the recent work AC25 , where the authors investigated how chirality introduced through a twist at the center of the particle, ’’transfer’’ into the macroscopic twist of heliconical $N_{TB}^{*}$ phase. Motivated by the results of this paper we studied the transfer of chirality from the local director chirality of the cholesteric into the chirality of the $N_{TB}^{*}$ phase. We show that the sign of the chiral heliconical $N_{TB}^{*}$ spiral should be always opposite to that of the $Ch$. Physics behind this relation is based on the orthogonality condition ${\bm{\varphi}}{\bm{n}}=0$ of the cholesteric director and vector order parameter of the $N_{TB}^{*}$ phase. To conclude, we note that the relationship between chirality signs at various incommensurate scales found in our work may be useful not only for the simple theoretical interpretation of the results presented in AC25 . One can have in mind, for example, the nontrivial dependencies on osmotic pressure of the optical dichroism signs of cholesteric liquid crystals formed by short DNA segments, discussed in the literature YS17 . Another seemingly distant and applied example concerns twistronic problems, VA22 - OV24 , where controlling the magnitude and sign of the chirality in meta-material layers is important.
Acknowledgments
This work has benefited from exchanges with many colleagues. I am strongly indebted to R.D.Kamien, who drew my attention to the effects of mutual influence of frustrated types of hexatic and nematic chiralities.
Список литературы
  • (1) L.E. Hough, M. Spannuth, M.Nakata, et al., ’’Chiral isotropic liquids from achiral molecules’’ Science, 325, 452 (2009).
  • (2) V.P.Panov, M.Nagaraj, J.K.Vij, et al., ’’Dynamic Calorimetry and XRD Studies of the Nematic and Twist-Bend Nematic Phase Transitions in a Series of Dimers with Increasing Spacer Length’’, Phys. Rev. Lett., 105, 167801 (2010).
  • (3) M.Cestari, S.Diez-Berart, D.A.Dunmur, et al., ’’Spontaneous polar and chiral symmetry breaking in ordered fluids - heliconical ferroelectric nematic phases’’, Phys. Rev. E, 84, 031704 (2011).
  • (4) V.Borshch, Y.K.Kim, J.Xiang, et al., ’’Nematic twist-bend phase with nanoscale modulation of molecular orientation’’, Nature Commun., 4, 2635 (2013).
  • (5) J.Mandle, E.J.Davis, S.A.Lobato, et al., ’’The relationship between molecular structure and the incidence of the $N_{TB}$ phase’’, Phys. Chem. Chem. Phys., 16, 6907 (2014).
  • (6) A. Ashkinazi, H.Chhabra, A.El.Moumane, M.M.C.Tortora, J.P.K. Doye, ’’Chirality transfer in lyotropic twist-bend nematics’’, arXiv:2508.03544v1 (2025).
  • (7) E.I. Kats, V.V. Lebedev, ’’Landau theory for helical nematic phase’’, JETP Lett., 100, 110 (2014).
  • (8) R.D. Kamien, ’’Liquids with Chiral Bond Order’’, J. Phys. II France, 6, 461 (1996).
  • (9) R.D. Kamien, D.R. Nelson, ’’Iterated Moire Maps and Braiding of Chiral Polymer Crystals’’, Phys. Rev. Letters, 74, 2499 (1995).
  • (10) R.D. Kamien, D.R. Nelson, ’’Defects in chiral columnar phases: Tilt-grain boundaries and iterated moire maps’’, Phys. Rev. E, 53, 650 (1996).
  • (11) R.D. Kamien, A.J. Levine, ’’Boundary Effects in Chiral Polymer Hexatics’’, Phys. Rev. Letters, 84, 3109 (2000).
  • (12) J.P. Straley, ’’Theory of piezoelectricity in nematic liquid crystals and of the cholesteric ordering’’, Phys. Rev. A 14, 1835 (1976).
  • (13) Y.M. Yevdokimov, S.G. Skuridin, S.V. Semenov, L.A. Dadinova, V.I. Salyanov, E.I. Kats, ’’Re-entrant cholesteric phase in DNA liquid-crystalline dispersion particles’’, J. Biol. Phys, 43, 45 (2017).
  • (14) C. Valagiannopoulos, ’’Electromagnetic Analog to Magic Angles in Twisted Bilayers of Two-Dimensional Media’’, Phys. Rev. Applied, 18, 044011 (2022).
  • (15) M. Orazbay, C. Valagiannopoulos, ’’ Twistronics-based polarization engineering’’, Phys. Rev. Applied, 22, 054028 (2024).